OpenStudy (anonymous):

For the pair of functions find the indcated domain f(x)=2x-5, g(x)=square root x+3 find the domain of f/g

OpenStudy (anonymous):

\[f(x)=2x-5, g(x)\sqrt{x+3}\]

OpenStudy (anonymous):

there is an = sign between the g(x) and the square root

OpenStudy (alfie):

It is asking the domain of \[\frac{f(x)}{g(x)} ?\]

OpenStudy (anonymous):

ya

OpenStudy (alfie):

\[\frac{2x-5}{\sqrt{x+3}}\] Conditions: \[\sqrt{x+3} \neq 0 \rightarrow x \neq-3\] \[x+3 \ge 0 \rightarrow x \ge -3\] Merging them both you get... \[x > -3\] So your function is defined in... \[(-3;+\infty)\]